OpenStudy (anonymous):

What is the value for x in this equation? 10x – 7 = 103x + 8

OpenStudy (anonymous):

alternate form -93-15=0 solution x=- 5/31

OpenStudy (anonymous):

103x-10x=93x -7-8=-15 x=-5/31